OpenStudy (anonymous):

Systems of Linear Inequalities Determine whether the ordered pair is a solution of the given system. A) (2.12); y>2x +4 y<3x +7 Solve each system of inequalities by graphing. B) y < 2x +4 -3x - 2y (greater than or equal too) 6

OpenStudy (anonymous):

Also remember to tag? @PeterPan ^.^ @TheKylieeM There's a way to determine A... just replace x with 2, and replace y with 12.

OpenStudy (anonymous):

Okay @PeterPan

OpenStudy (anonymous):

So, in both inequalities, y>2x+4 y<3x+7 Replace x with 2 and y with 12 See if it fits :)

OpenStudy (anonymous):

So @PeterPan the first one would be y>2(2)+4? then multiply the 2(2)?

OpenStudy (anonymous):

Don't forget to replace y with 12 :)

OpenStudy (anonymous):

Yes I know I just did that one which I got y<36 + 7 so my answer for both would be y>8 y<43

OpenStudy (anonymous):

@PeterPan right?

OpenStudy (anonymous):

y<43 ? How did you get? >.>

OpenStudy (anonymous):

I added the 7 and the 36 I got from y<3(12) + 7

OpenStudy (anonymous):

because 3(12) is 36... i think.

OpenStudy (anonymous):

but that was an x :/ Why did you replace it with 12? I said replace x with 2 replace y with 12 Careful next time, okay? ^.^

OpenStudy (anonymous):

Oh! Okay! Hold on! Again! lol

OpenStudy (anonymous):

Okay @PeterPan now I got for answers: 12>8 && 12 <10 so the first one should be the solution for the given set.

OpenStudy (anonymous):

12 < 10 ? Redo :)

OpenStudy (anonymous):

yes. because 3(2) is 6 and 6 + 4 = 10

OpenStudy (anonymous):

oh wait! I remember now!

OpenStudy (anonymous):

But it was y<3x +7 not y<3x +2

OpenStudy (anonymous):

oh I accidently put 4! instead of 7!

OpenStudy (anonymous):

so then is would be 12<13 @PeterPan

OpenStudy (anonymous):

Yeah, much better :) So... is 12 > 8 true?

OpenStudy (anonymous):

Yes that is true! So yes the ordered pair is a solution for equation A

OpenStudy (anonymous):

YES it is a solution for the FIRST INEQUALITY, but is it a solution for the second inequality? Is 12 < 13 true?

OpenStudy (anonymous):

Yes it is because 13 is greater than 12! :)

OpenStudy (anonymous):

and 12 is less than 13! :)

OpenStudy (anonymous):

Then, it's settled. (2,12) is a solution of the system :)

OpenStudy (anonymous):

Whoo hoo! Next question! :D

OpenStudy (anonymous):

hmmm... this is gonna be a tricky one :) Can you graph lines?

OpenStudy (anonymous):

Yes I can but I did look at this equation and I am VERY confused on how to graph something like this....

OpenStudy (anonymous):

Yes, don't worry :) First, pretend they're lines, and graph them... y = 2x +4 -3x - 2y = 6

OpenStudy (anonymous):

I'm still confused on how to graph lines like that... :|

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

Hmm... first put them in this form: y = mx + b

OpenStudy (anonymous):

and how do I do that exactly? Also what do each letter stand for?

OpenStudy (anonymous):

@PeterPan keep forgetting to tag you! ooops lol

OpenStudy (anonymous):

Only tag me once. What I want you to do is rearrange the equation so that y stands alone on one side of the equation :)

OpenStudy (anonymous):

like 2y - (-3x) (greater than or equal too) 6?

OpenStudy (anonymous):

and y is already standing alone on the first equation

OpenStudy (anonymous):

forget that they are inqualities, for now. -3x - 2y = 6

OpenStudy (anonymous):

okay but could it also be like the one I just gave but 2y - (-3x) = 6?

OpenStudy (anonymous):

No... because 2y - (-3x) = 2y+3x which isn't the same as -2y - 3x

OpenStudy (anonymous):

but you just flip the inequality sign right? dont you have to do that when theres a negative?

OpenStudy (anonymous):

Yeah, sure, but you sure failed to negate the right side (6)

OpenStudy (anonymous):

negate?

OpenStudy (anonymous):

make negative.

OpenStudy (anonymous):

Oh! ...................................

OpenStudy (anonymous):

so then how can i be able to make the y stand alone? I am really confused... o.0

OpenStudy (anonymous):

Hang on...

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

Here, an example :) \[\Large ax + by = c\]to isolate y, we bring all those without y to the right-side \[\Large by=-ax+c\]Then divide everything by b, the coefficient of y... \[\Large y = \frac{-ax+c}{b}\]That's how to isolate y :)

OpenStudy (anonymous):

Okay so to get why by its self I have to divide -3x - 2 by 6??

OpenStudy (anonymous):

y***

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

grrrrrrrrrrrrr......................

OpenStudy (anonymous):

Let's do it slowly :) \[-3x-2y=6\]Let's bring x to the right side \[-2y=3x+6\]

OpenStudy (anonymous):

Oh! I thought that the 2 had to be on the same side as the 3

OpenStudy (anonymous):

Not really, as long as you can get the y alone on one side, it should work :)

OpenStudy (anonymous):

okay so next step?

OpenStudy (anonymous):

Well, on the left side, it still says -2y, so we better divide everything by -2, to get rid of that coefficient, we get \[\Large y = \frac{3x-6}{-2}\]And there you go, y is alone on one side :)

OpenStudy (anonymous):

Oh! okay! so then?

OpenStudy (anonymous):

Well, simplify... \[\Large y = -\frac32x+3\]

OpenStudy (anonymous):

I'm going to graph this line, and you do the other, okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

The coordinate system :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

brb!

OpenStudy (anonymous):

The way to graph a line is to get its intercepts. Its y-intercepts and x-intercepts. Its x-intercept is obtained by letting y = 0, indeed, you can check this, if y = 0, then x = 2 So its x-intercept is the point (2,0)